AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:18:30 2025
Author
*
:
Subject
*
:
<p>22:30 Shift handed over. DAQ OK. Target OK. Beam OK.</p> <p> </p> <p>00:22 Temperatures OK</p> <p>Stats OK</p> <p>Good events OK</p> <p>Scalers OK</p> <p>System wide checks as usual - WR fails by 68 ns</p> <p>Leakage current OK</p> <p>Target OK</p> <p>Motors OK</p> <p>Pressure in ring 5E-10 mbar OK</p> <p>Beam OK</p> <p> </p> <p>02:26 - All OK as above</p> <p>WR now fails by 80</p> <p>Data rate ~800 KB/s</p> <p> </p> <p>03:40 All OK as above</p> <p>Data rate back to ~ 1900 KB/s</p> <p> </p> <p>04:20 All OK as above</p> <p>Started compressing files in WS3, monitor 1</p> <p>Noticed bias is not as stable as it has been so far, but no reason for concern. See attach.</p> <p> </p> <p>04:30 Saw a spike in counts similar to beam halo. See attachment. Did not see it before. No effects on leakage current. To be checked offline.</p> <p> </p> <p>06:00 Could not see any more spikes</p> <p>All other checks OK as before.</p> <p>Data rate back to ~1000 Kb/s</p> <p> </p> <p>06:32 Requested beam off for end of beamtime.</p> <p>Left DAQ on - run R88</p> <p>Turning off jet target as per Nikos instructions</p> <p>Final cleanup + pulse walkthrough to be done by JM and OH later today.</p> <p> </p> <p>Target nozzle removed - pressure in S1 now 1.93E-8 mbar</p> <p>Valves closed</p> <p>Entering cave to switch compressor off and change N backing pressure to 10 bar.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown